OpenStudy (anonymous):

The skybirdge of Petronas Twin Tower is 170m tall above the ground. A man on the bridge throws a ball vertically upwards at 8ms^-1 a) what is the velocity of the ball when it reaches it maximum height b)determine the velocity of the ball when it touches the ground c) find the time for the ball to reach the ground

OpenStudy (abmon98):

a)When any object is thrown upward in the direction away from gravity then it reaches to a certain height and again comes back to the floor with the same velocity it started with and the same take it took to reach its maximum height. At the Maxium Height Velocity Equals 0m/s b) 8m/s c) V^2=U^2+2as V=u+at -8=-10t t=-8/-10 0.8